OpenStudy (phagey):

The length of an object’s shadow varies directly with its height. If a can that is 13 cm tall casts a shadow that is 7.5 cm long, a bottle that is 25 cm tall will cast a shadow that is approximately 14.42 cm long. True False

OpenStudy (phagey):

@anglina4196 @mathmale

OpenStudy (phagey):

i think its false

OpenStudy (phagey):

wait so its true

OpenStudy (benji):

It's a ratio so if \[\frac{ 13 }{ 7.5 } = \frac{ 25 }{ x }\] So 13 is to 7.5 as 25 is to x. If x = 14.42, then the statement is true. we can start this by cross multiplying and we get\[13x = 25(7.5)\]Solve for x, what do you get?

OpenStudy (phagey):

14.42

OpenStudy (benji):

Nice :) So it's true

OpenStudy (phagey):

thanks @Benji and @anglina4196

OpenStudy (anglina4196):

your welcome @phagey